What is the y-value in the solution to the system {x+y=4}{2x-2y=0}
Answer by stanbon(75887)

What is the y-value in the solution to the system {x+y=4}{2x-2y=0}
------------------
x+y = 4
2x-2y = 0
----------------
Multiply thru 1st equation by 2 to get:
2x+2y = 8
2x-2y = 0
-----------------
Subtract 2nd eq. from 1st and solve for "y":
4y =8
y = 2
==============
Cheers,
Stan H.